OpenStudy (anonymous):

Confused help thank you. A mountain has a rise of 8,480 feet from its base. If there is a 2 degree drop in temperature for every rise of 530 feet in altitude, how many degrees difference is there between the ground temperature and the temperature at the peak?

OpenStudy (anonymous):

What's the mountain called

OpenStudy (anonymous):

help really confused

OpenStudy (anonymous):

(3,6)

OpenStudy (anonymous):

hmm somethin isnt right is there a different way?

OpenStudy (anonymous):

No, that;s the only way.

OpenStudy (anonymous):

yep ur right thanks for the help

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

8480 / 530 = 16 16 * 2 = 32 degree drop

OpenStudy (anonymous):

Any question, lala?

OpenStudy (anonymous):

yes thank you for helping me figure it out thank you

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

u deserve a medal

OpenStudy (anonymous):

I just like to help and it's my pleasure :)

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

Post whatever you need help, I'm here :)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

A square floor has an area of 361 square ft. What is the measure of one side of the floor?

OpenStudy (anonymous):

how do I find that out

OpenStudy (anonymous):

A = a^2 = 361 -> a = sqrt ( 361) = 19 ft

OpenStudy (anonymous):

oh okay I was doing it wrong thank you

OpenStudy (anonymous):

:)